OpenStudy (anonymous):

Can anyone tell me the first step on solving this? log1/2 (x-1) - log1/2 (x+1) < log1/2 (x-2) + 1

OpenStudy (anonymous):

first step is to try to combine in a single log, so you can write in equivalent exponential form

OpenStudy (anonymous):

lol they all have the same base so i can equal, right? i forgot that

OpenStudy (anonymous):

(x-1) - (x+1) < (x-2) + 1 can i do this?

OpenStudy (anonymous):

All I got was x<x^2

OpenStudy (anonymous):

lol i got x>-1 and with the conditions i got x>2 but thats not in my options

OpenStudy (anonymous):

thx anyways mads <3

OpenStudy (anonymous):

@KingGeorge any ideas?

OpenStudy (kinggeorge):

Can you clarify what the expression is in the equation editor?

OpenStudy (anonymous):

1/2 is the base. Ok, I will try

OpenStudy (kinggeorge):

1/2 is the base. Then, while you're idea is a good one, we can't do it yet because there's a term that isn't in a the same log function yet. Namely, the term 1.

OpenStudy (anonymous):

I thought of that. May I use the conversion of bases formula with that one?

OpenStudy (kinggeorge):

However, to put 1 in a function, note that \(log_{1\over2}(x)=1 \longrightarrow x= {1\over2}\) So instead of 1, use \[log_{1\over2}\left({1\over2}\right)\]Then you can take away the log functions because they're all equal.

OpenStudy (anonymous):

x= 3 or -1

OpenStudy (anonymous):

I redid the math on paper..

OpenStudy (anonymous):

\[\log _{1/2} (x-1) - \log _{1/2} (x+1) < \log _{1/2} (x-2) + 1\]

OpenStudy (anonymous):

oh wait greater than sign...

OpenStudy (kinggeorge):

So \((x-1)-(x+1)<(x-2)+1/2\) so \(-2 < x-3/2\) so we finally get that \[x>-{1\over2}\]

OpenStudy (anonymous):

That's weird... I will post my options

OpenStudy (anonymous):

I made a mistake with my second calculation too lol!

OpenStudy (kinggeorge):

Nevermind, I made a mistake too. You can only remove the logs if the entire side is contained within a single logarithm. So instead, put the log around the 1, combine the logs, and then remove them.

OpenStudy (anonymous):

a) \[\left\{ x R / 2 \le x \le 3\right\}\] b) \[\left\{ x \in R /2<x <3 \right\}\] c) \[\left\{ xR/0<x <3\right\}\] d) \[\left\{ xR/0 \le x \le3 \right\}\]

OpenStudy (anonymous):